OpenStudy (anonymous):

Suppose that to send a telegram it costs $2 for the first 10 words in the telegram and Y cents for each additional words. a. Write an equation to find the cost C of a telegram 26 words long. b.Solve this equation for Y. PLZ show me how to do this.

OpenStudy (anonymous):

a) C = 2 + 16 x (Y/100) (where C is in dollars) b) Rearrange this: C - 2 = 16 x (Y/100) which can be written as: Y = (C - 2) x 100/16

OpenStudy (anonymous):

I am giving a generalized solution below and solving for given specific condition C= 2+ (Y/100) * (x-10) dollars. Where x is the number of words here a) C = 2+ (Y/100)*(16) $ = (200+16Y)/(100) b) a => Y= (100C-200)/16